# Quantum-enhanced mean value estimation via adaptive measurement

### v3

@@ Abstract
Quantum-enhanced (i.e., less query complexity compared to any classical method) mean value estimation of observables is a fundamental task in various quantum technologies; in particular, it is an essential subroutine in quantum computing algorithms. Notably, the quantum estimation theory identifies the ultimate precision of such estimator, which is referred to as the quantum Cramér-Rao (QCR) lower bound or equivalently the inverse of the quantum Fisher information. Because the estimation precision directly determines the performance of those quantum technological systems, it is highly demanded to develop a generic and practically implementable estimation method that achieves the QCR bound. Under imperfect conditions, however, such an ultimate and implementable estimator for quantum mean values has not been developed. In this paper, we propose a quantum-enhanced mean value estimation method in a depolarizing noisy environment that asymptotically achieves the QCR bound in the limit of a large number of qubits. To approach the QCR bound in a practical setting, the method adaptively optimizes the amplitude amplification and a specific measurement that can be implemented without any knowledge of state preparation. We provide a rigorous analysis for the statistical properties of the proposed adaptive estimator such as consistency and asymptotic normality. Furthermore, several numerical simulations are provided to demonstrate the effectiveness of the method, particularly showing that the estimator needs only a modest number of measurements to almost saturate the QCR bound.
@@ 1 Introduction
For this purpose, the statistical estimation theory [21, 22] and their quantum extension [23, 24] are useful. In particular, the seminal Cramér-Rao inequality and Fisher information, which identify the limit of estimation precision, have been extended to quantum versions [25, 26, 24, 27, 28]; that is, the quantum Cramér-Rao (QCR) inequality and quantum Fisher information. Actually, as in the classical case, they can be used to determine the ultimate performance of several quantum information processing systems such as quantum sensors [29, 30, 31, 32]. Moreover, the QCR acts as a fundamental limitation on information extractable from quantum mechanical systems, e.g., [33]. Therefore, extensive investigations have been conducted to characterize estimators that (almost) achieve the ultimate precision limit i.e., the QCR bound [34, 35, 36, 37, 38, 39, 40, 41]. Note that whether such optimal estimators are implementable depends on available quantum operations in the target system. Devising an implementable and optimal (in the sense of QCR) estimator is thus one of the most important subjects in quantum sensing and metrology.
@@
In this way, the quantum statistical estimation theory should be appropriately applied to develop efficient estimators for various quantities in quantum computing, such as mean values of quantum observables. In particular, toward the early success of quantum computing, there is an increasing need for efficient estimators that can be realized with as few demanding quantum operations as possible. This exploration has only recently started. For instance, some less-demanding implementation methods for the quantum-enhanced (i.e., less query complexity compared to any classical means) phase estimation algorithm [42, 43, 44, 45] and amplitude estimation algorithm [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56] have been developed, and their performance has been investigated in depth using statistical estimation theory, yet without studying the QCR bound. Moreover, Ref. [18] formulated the mean-value estimation problem within the framework of amplitude estimation and proposed a Bayesian inference method for this problem; however, again, there was no discussion on how to achieve the QCR bound. To our best knowledge, only Ref. [53] gave a detailed analysis on the QCR bound in the quantum amplitude estimation problem; in particular, under the assumption that the system is subjected to depolarization noise, the estimator given in [53] achieves the precision that is yet strictly lower than the ultimate limit.
@@
Therefore, in a broad sense, it is important to explore the ultimate and implementable quantum-enhanced estimator for a particular estimation problem under a noisy environment. In this paper, we focus on the problem of constructing a less-demanding quantum mean-value estimator in a depolarizing noisy environment. In light of quantum estimation theory and quantum metrology, the following questions are important. (i) What is the quantum-enhanced precision limit (i.e., the QCR bound) for estimating a quantum mean value in a less-demanding way? (ii) Can we construct an implementable estimator such that the estimation precision achieves the QCR bound? (iii) Also, is there any mathematical guarantee of the estimation performance? The main contribution of this paper is to give answers to all the above questions.
@@
A summary of those contributions are as follows. The details are provided in Section 3. First, the answer to the question (i) is given in Theorem 1; we derive the precision limit for estimating a quantum mean value, in the setup where we are allowed to use noisy quantum states but no demanding controlled amplifications. Based on the findings obtained from the precision limit, we construct a modified maximum likelihood estimator that adaptively optimizes a specific measurement (positive operator-valued measure (POVM)) and the number of amplitude amplification operations, by maximizing the magnitude of quantum enhancement. The optimality of this estimator is supported by Theorem 2, stating that there exists a POVM that makes the ratio of the quantum and classical Fisher information close enough to 1 (leading that the QCR bound is almost achieved) in the limit of large number of qubits. Note that, as mentioned in the second paragraph, the implementation of such an optimal estimator is in general nontrivial, but our estimator can be efficiently implemented without any knowledge of the target state, which is usually prepared by a black-box unitary in the mean value estimation problem. Thus, this is the answer to the question (ii). As for the question (iii), we actually prove that the estimator satisfies the following statistical properties.
@@ 2 Preliminary
In the second step, we estimate the amplitude based on the measurement results obtained in the first step by the maximum likelihood estimation method. Note that in the first step, we prepare the quantum states with different resolutions in order to eliminate the ambiguities, as similar to the methods for the optical phase measurement [57, 58, 59, 60]. Specifically, for each of predefined odd numbers $2m_{k}+1~{}(k=1,2,\cdots,M)$, we prepare the state (2) and measure it $N$ times with the computational basis independently. Here, $M$ denotes the total number of types of quantum states to be prepared. We write $x^{(k)}\in\{0,1,\cdots,N\}$ as the number of hitting "1" for the state $\mathcal{Q}^{m_{k}}\mathcal{A}\ket{0}_{n+1}$. The measurement results for $M$ different states are put together as $\bm{x}_{M}:=(x^{(1)},x^{(2)},\cdots,x^{(M)})$. Then the likelihood function to have $\bm{x}_{M}$ is given by
@@ 3 Quantum-enhanced mean value estimation
First, in Section 3.1, we formulate the quantum mean value estimation problem in a similar way to the amplitude estimation problem shown in the previous section. Then, we provide the precision limit in estimating the mean value when we can prepare and measure noisy amplitude-amplified quantum states, in Section 3.2. To achieve the precision limit, we next provide the efficient and implementable measurements in Section 3.3 and thereby develop an adaptive mean-value estimation algorithm in Section 3.4. In Section 3.5, we discuss the statistical properties of our method such as consistency, asymptotic normality, and the (classical) Fisher information that can become large by the adaptive optimization.
@@ 3.1 Mean value estimation problem
Our goal is to estimate the mean value of a known Hermitian operator $\mathcal{O}$ with eigenvalues ±$\pm 1$, i.e., $\langle\mathcal{O}\rangle:=\bra{A}\mathcal{O}\ket{A}$, where |$\ket{A}:=A\ket{0}_{n}$ is an $n$-qubit quantum state with a state preparation unitary $A$. Here, we assume that $A$ and $A^{\dagger}$ are black-box operations yet implementable on a device, similar to $\mathcal{A}$ in the previous section. Note that $\mathcal{O}$ is a unitary operator, which is assumed to be implementable; also, the (Ludër’s) projective measurement for $\mathcal{O}$ is implementable, which is often assumed in quantum algorithms. A typical $\mathcal{O}$ is a single Pauli string such as $\mathcal{O}=X\otimes Z\otimes\cdots$; note that our method is applicable to estimate the mean value of a linear combination of Pauli strings that form a completely anti-commuting set, because there exists a unitary operator that groups the linear combination into a single Pauli string [61].
@@
Recall that the oracles ${A}$ and ${A}^{\dagger}$ are implementable, meaning that the amplification operator ${Q}$ is also implementable. This operator keeps the subspace $\mathcal{S}$ invariant. Similar to Eq. (3), the representation of $Q$ on $\mathcal{S}$ is expressed as $\left.Q\right|_{\mathcal{S}}={\rm exp}(i\theta^{*}\bar{Y})$, where we define $\bar{Y},\bar{Z}$ as the Pauli Y, Z matrices in the basis |$\ket{\bar{0}}$ and |$\ket{\bar{1}}$. The rotational angle $\theta^{*}$ is related to the target mean value as $\theta^{*}:=\arccos\left(\langle\mathcal{O}\rangle\right)$; in what follows $\theta^{*}$ is also referred to as the target value. Note that, for the target mean value $\langle\mathcal{O}\rangle=\cos\theta^{*}$, the domain of $\theta^{*}$ is given by $\theta^{*}\in(0,\pi)$, which is twice as that of $\phi^{*}$ in the previous section.
@@
where this state uses $2(m+1)$ queries to $A$ and $A^{\dagger}$ in total. Here, the basis $A^{\dagger}\ket{\bar{1}}$ is actually unknown by the black-box assumption of $A^{\dagger}$; thus, the class of implementable measurements on the state (13) is restricted, as detailed in the following sections. On the other hand, the projective measurement of $\mathcal{O}$ on $Q^{m}\ket{\bar{0}}$ should be considered in $\mathcal{S}$ because of the implicit $\theta^{*}$-dependence of $\mathcal{O}$. That is, this process is equivalent to the projective measurement of $\bar{Z}\oplus\left.\mathcal{O}\right|_{\mathcal{S}^{\perp}}$ on the following state with $12m+$ 1 queries to $A$ and $A^{\dagger}$:
@@
This means that we can measure the quantum state with $(2m+1)$-fold enhanced resolution of $\theta^{*}$, although available POVM is restricted. Therefore, for a given $\alpha\in\mathbb{N}$, we can prepare and measure the quantum state Eq. (13) ($\alpha=2m+2$) or Eq. (14) ($\alpha=2m+1$), according to the parity of $\alpha$; we write these states in a unified manner as $\rho(\alpha;\theta^{*})$. In the following, we call $\alpha$ the amplified level. Note that $\rho(\alpha;\theta^{*})$ requires $\alpha$ queries to $A$ and $A^{\dagger}$ in total.
@@ 3.2 Limit of estimation precision with noisy amplitude-amplified states
Here, we assume the state $\rho(\alpha;\theta^{*})$ is subjected to the (global) depolarization noise with probability $1-p_{\rm q}~{}(>0)$ for each use of $A$ ($A^{\dagger}$), as in the previous case. Then, the output state with $\alpha$ queries to $A$ and $A^{\dagger}$ is given by
@@
For a given number $N_{\rm q}$ of queries to $A$ and $A^{\dagger}$, let us prepare a set of $\rho_{\mathcal{D}}(\alpha;\theta^{*})$ by splitting the total queries. More specifically, we consider any partition $\{\alpha^{\prime}_{k}\}$ of a given total queries $N_{\rm q}$ such that $N_{\rm q}=\sum_{k=1}^{M^{\prime}}\alpha^{\prime}_{k}$ for some positive integer $M^{\prime}$. Then, we define the quantum state before measurement as
@@
Note that each $\rho_{\mathcal{D}}(\alpha^{\prime}_{k};\theta^{*})$ uses $\alpha^{\prime}_{k}$ queries to $A$ and $A^{\dagger}$, and thus the quantum state $\rho_{\mathcal{D}}^{{\rm(bm)}}$ uses $N_{\rm q}$ queries in total. Here, for all partitions, we provide the upper bound of the total quantum Fisher information $\mathcal{I}_{\rm q,tot}[\rho_{\mathcal{D}}^{{\rm(bm)}}]$ of the state $\rho_{\mathcal{D}}^{{\rm(bm)}}$ in estimating $\theta^{*}$.
@@
The proof is given in Appendix B, and a similar evaluation of quantum Fisher information can also be found in the previous work on noisy quantum metrology [66]. Note that, if the argmax returns multiple values, we take the minimum as $\alpha_{\rm B}$. A particularly important fact is that the equality in the case $N_{\rm q}=r\alpha_{\rm B}>\alpha_{\rm B}$ holds for the partition $M^{\prime}=r$ and $\alpha^{\prime}_{k}=\alpha_{\rm B}~{}\forall k$; that is,
@@ 3.3 The two-type measurement
where the superscript even means that we can perform this measurement on the quantum state $\rho_{\mathcal{D}}$ with even $\alpha$ (i.e., $\alpha=2m+2$). Actually, this POVM can be implemented without any knowledge of the black-box operations $A$ and $A^{\dagger}$, by the computational basis measurement followed by post-processing for classifying the obtained measurements into "0" or "1".
@@
Here $\alpha\in\mathbb{N}$ represents the number of queries to the operator $A$ and $A^{\dagger}$. We perform the measurement via Eq. (3.3) when $\alpha$ is odd $(\alpha=2m+1)$ and via Eq. (3.3) when $\alpha$ is even $(\alpha=2m+2)$. Also, the classical Fisher information corresponding to each measurement is represented as

### v4

@@ Abstract
Quantum-enhanced (i.e., higher performance by quantum effects than any classical methods) mean value estimation of observables is a fundamental task in various quantum technologies; in particular, it is an essential subroutine in quantum computing algorithms. Notably, the quantum estimation theory identifies the ultimate precision of such an estimator, which is referred to as the quantum Cramér-Rao (QCR) lower bound or equivalently the inverse of the quantum Fisher information. Because the estimation precision directly determines the performance of those quantum technological systems, it is highly demanded to develop a generic and practically implementable estimation method that achieves the QCR bound. Under imperfect conditions, however, such an ultimate and implementable estimator for quantum mean values has not been developed. In this paper, we propose a quantum-enhanced mean value estimation method in a depolarizing noisy environment that asymptotically achieves the QCR bound in the limit of a large number of qubits. To approach the QCR bound in a practical setting, the method adaptively optimizes the amplitude amplification and a specific measurement that can be implemented without any knowledge of state preparation. We provide a rigorous analysis for the statistical properties of the proposed adaptive estimator such as consistency and asymptotic normality. Furthermore, several numerical simulations are provided to demonstrate the effectiveness of the method, particularly showing that the estimator needs only a modest number of measurements to almost saturate the QCR bound.
@@ 1 Introduction
For this purpose, the statistical estimation theory [21, 22] and their quantum extension [23, 24] are useful. In particular, the seminal Cramér-Rao inequality and Fisher information, which identify the limit of estimation precision, have been extended to quantum versions [25, 26, 24, 27, 28]; that is, the quantum Cramér-Rao (QCR) inequality and quantum Fisher information. As in the classical case, they can be used to determine the ultimate performance of several quantum information processing systems such as quantum sensors [29, 30, 31, 32, 33, 34]. Moreover, the QCR acts as a fundamental limitation on information extractable from quantum mechanical systems, e.g., [35]. Therefore, extensive investigations have been conducted to characterize estimators that (almost) achieve the QCR lower bound i.e., the ultimate precision bound for optimized quantum measurement [36, 37, 38, 39, 40, 41, 42, 43, 44]. Note that whether such optimal estimators are implementable depends on available quantum operations in the target system. Devising an implementable and optimal (in the sense of QCR) estimator is thus one of the most important subjects in quantum sensing and metrology.
@@
In this way, the quantum statistical estimation theory should be appropriately applied to develop efficient estimators for various quantities in quantum computing, such as mean values of quantum observables. At the same time, toward the early success of quantum computing, there is an increasing need for efficient estimators that can be realized with as few demanding quantum operations as possible. For instance, some less-demanding implementation methods for phase estimation algorithm [45, 46, 47, 48, 49, 50] and amplitude estimation algorithm [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] have been developed, and their performance has been investigated using the statistical estimation theories. Nevertheless, in practical noisy settings before the fault-tolerant era, devising an estimator that is both easy to implement and capable of achieving the ultimate QCR bound remains an ongoing subject of research. Focusing on the mean-value estimation problem in a noisy environment, Ref. [18] formulated this problem within the framework of amplitude estimation and proposed a Bayesian inference method for this problem; however, there was no discussion on how to achieve the QCR bound. To our best knowledge, only Ref. [58] gave a detailed analysis on the QCR bound in the noisy amplitude estimation problem; in particular, under the assumption that the system is subjected to depolarization noise, the estimator given in [58] achieves the precision that is yet strictly lower than the ultimate limit.
@@
In this paper, we focus on the problem of constructing a less-demanding quantum mean-value estimator in a depolarizing noisy environment. In light of quantum estimation theory and quantum metrology, the following questions are important. (i) What is the quantum-enhanced precision limit (or, the optimal QCR bound) over less-demanding protocols for estimating a quantum mean value? (ii) Can we construct an implementable estimator such that the estimation precision achieves the QCR bound? (iii) Also, is there any mathematical guarantee of the estimation performance? The main contribution of this paper is to give answers to all the above questions.
@@
A summary of those contributions are as follows. The details are provided in Section 3. First, the answer to the question (i) is given in Theorem 1; we derive the precision limit for estimating a quantum mean value, in the setup where we are allowed to use noisy quantum states but no demanding controlled amplifications. Based on the findings obtained from the precision limit, we construct a modified maximum likelihood estimator that adaptively optimizes a specific measurement (positive operator-valued measure (POVM)) and the number of amplitude amplification operations, by maximizing the magnitude of quantum enhancement. The optimality of this estimator is supported by Theorem 2, stating that there exists a POVM that makes the ratio of the quantum and classical Fisher information close enough to 1 (leading that the QCR bound is almost achieved) in the limit of large number of qubits. Note that, as mentioned in the second paragraph, the implementation of such an optimal estimator is in general nontrivial, but our estimator can be efficiently implemented without any knowledge of the target state, which is usually prepared by a black-box unitary in the mean value estimation problem. Thus, this is the answer to the question (ii). As for the question (iii), we prove that the estimator satisfies the following statistical properties.
@@ 2 Preliminary
In the second step, we estimate the amplitude based on the measurement results obtained in the first step by the maximum likelihood estimation method. Note that in the first step, we prepare the quantum states with different resolutions in order to eliminate the ambiguities, as similar to the methods for phase measurement [62, 46, 63, 64, 65, 50]. Specifically, for each of predefined odd numbers $2m_{k}+1~{}(k=1,2,\cdots,M)$, we prepare the state (2) and measure it $N$ times with the computational basis independently. Here, $M$ denotes the total number of types of quantum states to be prepared. We write $x^{(k)}\in\{0,1,\cdots,N\}$ as the number of hitting "1" for the state $\mathcal{Q}^{m_{k}}\mathcal{A}\ket{0}_{n+1}$. The measurement results for $M$ different states are put together as $\bm{x}_{M}:=(x^{(1)},x^{(2)},\cdots,x^{(M)})$. Then the likelihood function to have $\bm{x}_{M}$ is given by
@@ 3 Quantum-enhanced mean value estimation
First, in Section 3.1, we formulate the quantum mean value estimation problem in a similar way to the amplitude estimation problem shown in the previous section. Then, we provide the precision limit in estimating the mean value when we can prepare and measure noisy amplitude-amplified quantum states, in Section 3.2. To achieve the precision limit, we next provide the efficient and implementable measurements in Section 3.3 and thereby develop an adaptive mean-value estimation algorithm in Section 3.4. In Section 3.5, we discuss the statistical properties of our method such as consistency, asymptotic normality, and the (classical) Fisher information that can become large by the adaptive optimization. Finally, a summary of our protocol is provided in Section 3.6.
@@ 3.1 Mean value estimation problem
Our goal is to estimate the mean value of a known Hermitian operator $\mathcal{O}$ with eigenvalues ±$\pm 1$, i.e., $\langle\mathcal{O}\rangle:=\bra{A}\mathcal{O}\ket{A}$, where |$\ket{A}:=A\ket{0}_{n}$ is an $n$-qubit quantum state with a state preparation unitary $A$. Here, we assume that $A$ and $A^{\dagger}$ are black-box operations yet implementable on a device, similar to $\mathcal{A}$ in the previous section. Note that $\mathcal{O}$ is a unitary operator, which is assumed to be implementable; also, the projective measurement $\{(I_{n}+\mathcal{O})/2,(I_{n}-\mathcal{O})/2\}$ is implementable, which is often assumed in quantum algorithms. A typical $\mathcal{O}$ is a single Pauli string such as $\mathcal{O}=X\otimes Z\otimes\cdots$; note that our method is applicable to estimate the mean value of a linear combination of Pauli strings that form a completely anti-commuting set, because there exists a unitary operator that groups the linear combination into a single Pauli string [66].
@@
[…]
@@
where this state uses $2(m+1)$ queries to $A$ or $A^{\dagger}$ in total. Here, the basis $A^{\dagger}\ket{\bar{1}}$ is actually unknown by the black-box assumption of $A^{\dagger}$; thus, the class of implementable measurements on the state (13) is restricted, as detailed in the following sections. On the other hand, the projective measurement of $\mathcal{O}$ on $Q^{m}\ket{\bar{0}}$ should be considered in $\mathcal{S}$ because of the implicit $\theta^{*}$-dependence of $\mathcal{O}$. That is, this process is equivalent to the projective measurement of $\bar{Z}\oplus\left.\mathcal{O}\right|_{\mathcal{S}^{\perp}}$ on the following state with $12m+$ 1 queries to $A$ or $A^{\dagger}$:
@@
This means that we can measure the quantum state with $(2m+1)$-fold enhanced resolution of $\theta^{*}$, although available POVM is restricted. Therefore, for a given $\alpha\in\mathbb{N}$, we can prepare and measure the quantum state Eq. (13) ($\alpha=2m+2$) or Eq. (14) ($\alpha=2m+1$), according to the parity of $\alpha$; we write these states in a unified manner as $\rho(\alpha;\theta^{*})$. In the following, we call $\alpha$ the amplified level. Note that $\rho(\alpha;\theta^{*})$ requires $\alpha$ queries to $A$ or $A^{\dagger}$ in total.
Here, we assume the state $\rho(\alpha;\theta^{*})$ is subjected to the (global) depolarization noise with probability $1-p_{\rm q}~{}(>0)$ for each use of $A$ ($A^{\dagger}$), as in the previous case. Then, the output state with $\alpha$ queries to $A$ or $A^{\dagger}$ is given by
We remark that $A$ and $A^{\dagger}$ (e.g., the time evolution of a global Hamiltonian) are usually comprised of a large number of gates, which would sufficiently scramble local noise into global depolarization noise. Such assumptions are verified in terms of theoretical [71, 72] and (small-size) experimental [57] aspects, while we leave further verification in a large-size experiment as future work.
@@ 3.2 Limit of estimation precision with noisy amplitude-amplified states
For a given number $N_{\rm q}$ of queries to $A$ or $A^{\dagger}$, let us prepare a set of $\rho_{\mathcal{D}}(\alpha;\theta^{*})$ by splitting the total queries. More specifically, we consider any partition $\{\alpha^{\prime}_{k}\}$ of given total queries $N_{\rm q}$ such that $N_{\rm q}=\sum_{k=1}^{M^{\prime}}\alpha^{\prime}_{k}$ for some positive integer $M^{\prime}$. Then, we define the quantum state before measurement as
@@
Note that each $\rho_{\mathcal{D}}(\alpha^{\prime}_{k};\theta^{*})$ uses $\alpha^{\prime}_{k}$ queries to $A$ or $A^{\dagger}$, and thus the quantum state $\rho_{\mathcal{D}}^{{\rm(bm)}}$ uses $N_{\rm q}$ queries in total. Here, for all partitions, we provide the upper bound of the total quantum Fisher information $\mathcal{I}_{\rm q,tot}[\rho_{\mathcal{D}}^{{\rm(bm)}}]$ of the state $\rho_{\mathcal{D}}^{{\rm(bm)}}$ in estimating $\theta^{*}$.
@@
The proof is given in Appendix B, and a similar evaluation (i.e., the loss of quadratic term on ${N_{\rm q}}$) of quantum Fisher information can also be found in the previous works on noisy quantum metrology [73, 74, 75]. Note that, if the argmax returns multiple values, we take the minimum as $\alpha_{\rm B}$. A particularly important fact is that the equality in the case $N_{\rm q}=r\alpha_{\rm B}>\alpha_{\rm B}$ holds for the partition $M^{\prime}=r$ and $\alpha^{\prime}_{k}=\alpha_{\rm B}~{}\forall k$; that is,
@@ 3.3 The two-type measurement
where the superscript even means that we can perform this measurement on the quantum state $\rho_{\mathcal{D}}$ with even $\alpha$ (i.e., $\alpha=2m+2$). This POVM can be implemented without any knowledge of the black-box operations $A$ and $A^{\dagger}$, by the computational basis measurement followed by post-processing for classifying the obtained measurements into "0" or "1".
@@
Here $\alpha\in\mathbb{N}$ represents the number of queries to the operator $A$ or $A^{\dagger}$. We perform the measurement via Eq. (3.3) when $\alpha$ is odd $(\alpha=2m+1)$ and via Eq. (3.3) when $\alpha$ is even $(\alpha=2m+2)$. Also, the classical Fisher information corresponding to each measurement is represented as